The cost in dollars of producing x headphones is

C(x) = 2000 + 150x – 0.1x2


a. Calculate the average cost of producing 250 headphones.


b. Calculate the marginal cost of producing the 250th headphone. 


1
Expert's answer
2021-03-25T01:52:23-0400

"a) C(x) = 2000+150x-0.1x^2\\
C(250) = 2000+150(250)-0.1(250)^2 = 2000+37500-6250=33250\\"

Average cost of x quantity = "\frac{C(x)}{x}"

So average cost of 250 headphones = "\frac{C(250)}{250} = \frac{33250}{250} = 133"

i.e The average cost of producing 250 headphones is 133 dollars.

b) Let M denote marginal cost.

We have that "M = \dfrac{dC(x)}{dx} = 150-0.2x"

"\therefore M(250) = 150-0.2(250) = 150- 50 = 100"

i.e The marginal cost of producing the 250th headphone is 100 dollars.